OpenStudy (anonymous):

The perimeter of a rectanlge is 22 ft. The length is 7 ft longer than the width. Find the dimenisons. let x be length and y be width.

OpenStudy (anonymous):

so 2x + 2y = 22 x = 7 + y so 2(7 + y) + 2y = 22 14 + 4y = 22 y = 2 so x = 9

OpenStudy (anonymous):

kool that is the way that i set it up but was worried thanks.

OpenStudy (anonymous):

P = 2L + 2W 22=2(x+7)+2x 22=2x+14+2x x=2 therefore L=9, W=2

OpenStudy (radar):

Let x = length w = width x=w+7 (given) 2(w+7)+2w=22 (the sum of the four sides is the perimeter. 2w+14+2w=22 4w=22-14=8 w=2 x=w+7 (given) x=9 so length =9 and width =2, Perimeter = 22, Length is 7 ft longer than width. It all fits.